OpenStudy (mynameisnemo):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 19

OpenStudy (mynameisnemo):

@agent0smith

OpenStudy (mynameisnemo):

@mjdennis @.Sam. last question then finals :)

sam (.sam.):

If you're back, take a look at this and I'll explain it further

OpenStudy (mynameisnemo):

one second let me read it

OpenStudy (mynameisnemo):

okay i read it

OpenStudy (mynameisnemo):

I'm stuck on that complex step

sam (.sam.):

\[(x^2 + 2x+?) + (y^2 + 4y+?) = 19 \] So, the half of 2 is 1, the squared of it is 1, as for 'y' the half of it is 2, squared it you get 4, add these to your RHS as well \[(x^2 + 2x+1) + (y^2 + 4y+4) = 19 + 1+4 \] \[(x^2 + 2x+1) + (y^2 + 4y+4) =24 \]

OpenStudy (mynameisnemo):

what what

OpenStudy (mynameisnemo):

wait what*

sam (.sam.):

lol, right, from here you'll have to find the question mark and you'll take the 'Number' from 2x, and 4y. then, there's a method you'll need to make, take the half of 2x and 4y and square it, but without the x and y. So, for x, half of 2 is 1, square of it is still 1. For you, half of 4 is 2, take the square of it is 4 again. Then you'd put 1 and 4 accordingly \[(x^2 + 2x+1) + (y^2 + 4y+4) = 19 + 1+4\]

OpenStudy (mynameisnemo):

ok where does the 1 and 4 go in the parenthesis?

sam (.sam.):

I mean For 'y', not 'For you'

sam (.sam.):

Put them as a number inside the parenthesis, with \((x^2+2x)\)

sam (.sam.):

You'd get \[(x^2 + 2x+1) + (y^2 + 4y+4) = 19 + 1+4\]

OpenStudy (mynameisnemo):

yes but after they're in there where do they go?

sam (.sam.):

Let them be there because you'll need to factorize the two equations later

OpenStudy (mynameisnemo):

okay, I have that written down, now what?

sam (.sam.):

Great now factorize x^2+2x+1

OpenStudy (mynameisnemo):

what do you mean by that?

sam (.sam.):

Factorize into this I mean \[(x+1)^2\]

sam (.sam.):

And you'll need to factorize y2+4y+4 as well

OpenStudy (mynameisnemo):

so \[(y+4)^2\]

OpenStudy (mynameisnemo):

?

sam (.sam.):

Nope your values are off, if you expand it you won't get y2+4y+4

sam (.sam.):

Try (y+2)^2

OpenStudy (mynameisnemo):

oh bc 2 is the square root of 4?

OpenStudy (mynameisnemo):

then I would square root the whole left side right? so it would look like \[(x+1) + (y + 2) = 24\]

sam (.sam.):

No just normal factorizing equation